OpenStudy (anonymous):

help evaluate 80*2^-0/28= 80*2^28/28= 80*2^56/28=

OpenStudy (anonymous):

\[\LARGE 80*\left(2\right)^{-0/28}=\] is this what you mean?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i got 160

OpenStudy (anonymous):

\[\LARGE \frac0b=0\] any number instead of b \[\Large b\neq0\] is equal to 0 so you have ...?

OpenStudy (unklerhaukus):

80

OpenStudy (anonymous):

\[\LARGE x^{-a}=\frac{1}{x^a}\]

OpenStudy (anonymous):

i really dnt get this...

OpenStudy (anonymous):

example?

OpenStudy (anonymous):

\[\LARGE 80\cdot \left(2\right)^{\frac{-0}{28}}=80\cdot 2^0=80\cdot 1\]

OpenStudy (anonymous):

oooh i got the steps right but instead i did 80*2 then i did 0 which gave me 160 thnks @Kreshnik

OpenStudy (anonymous):

you're welcome ... can you do the rest ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

is expression 80*2^2

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

\[\LARGE 80\cdot \left(2\right)^{\frac{28}{28}}= 80\cdot 2^1\]

OpenStudy (anonymous):

yeah just got the mistake

OpenStudy (anonymous):

so 160

OpenStudy (anonymous):

320 is the last one

OpenStudy (anonymous):

What does each value of the expression represent?

OpenStudy (anonymous):

how do you mean represent?

OpenStudy (anonymous):

nvm i got it

OpenStudy (phi):

@daja2fly Remember exponents are negative. You are modeling radioactive decay, and the amount of a substance DECAYS (gets smaller) with time. So for example, 2^-2 is 1/4 so \[ 80\cdot 2^{-2} = 80\cdot \frac{1}{4}= 20 \] Notice that you have less of the substance. Btw, when you evaluate an expression you do exponents before multiplication.

OpenStudy (phi):

If you have a positive exponent, you have exponential GROWTH. That is what populations do, for example.

OpenStudy (anonymous):

so will the third be -40

OpenStudy (phi):

Negative exponents work differently. You have to learn new rules (which may not make much sense, but they work) So for the third question \( 80 \cdot 2^{-\frac{56}{28}} \) You can evaluate (simplify) -56/28 to -2 So you have \( 80 \cdot 2^{-2} \) Now use the rule that a NEGATIVE EXPONENT means flip: \[ 2^{-2}= \frac{1}{2^{2}} \] and \[ 80 \cdot \frac{1}{4} \] This is not -40

OpenStudy (anonymous):

the answers r 80, 40 and 20

OpenStudy (phi):

Yes, do you understand how to get them? The first means after 0 years (no time has gone by) you have what you started with after 1 year, you have half of what you started with (that is why they call it "half-life")

OpenStudy (anonymous):

yup